Acolgrove:

The ratio of roosters to chickens is 2 to 8. If there are 5 roosters, how many chickens are there? Complete the ratio table to solve. Roosters 2 1 5 Chickens 8 10 12 18 20

Maple:

So it should also be 50% Rooster. So if it is \(5\) It would be \(5\times2=10\)

Maple:

Instead of also, It should say always*

Acolgrove:

oaky tysm

Acolgrove:

okay(*

ramen:

That is incorrect tho.

Acolgrove:

oop

ramen:

Since the ratio is 2:8 that means we can simplify it to 1 rooster for every 4 chickens. For 5 roosters, since 1*5=5 roosters, 4*5=20 chickens so for 5 roosters there are 20 chickens not 10.

Maple:

True, my bad. Ramen=Big Brain I=Small brain.

Acolgrove:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Maple True, my bad. Ramen=Big Brain I=Small brain. \(\color{#0cbb34}{\text{End of Quote}}\) lol

darkknight:

this: So it should also be 50% Rooster. is an incorrect statement Ratio is 2:8

Acolgrove:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ramen Since the ratio is 2:8 that means we can simplify it to 1 rooster for every 4 chickens. For 5 roosters, since 1*5=5 roosters, 4*5=20 chickens so for 5 roosters there are 20 chickens not 10. \(\color{#0cbb34}{\text{End of Quote}}\) okay tysm Ramen

ramen:

Np

ramen:

Btw the ratio isn't 50%. If it was 50% it would be 4:8 not 2:8

Maple:

My mistake, That was a big mistake. I know it's 25%.

ramen:

It's ight